The perimeter of a rectangle yard is 290 ft. If its length is 25 feet greater than its width, what are the dimensions of the yard?
Answer by checkley77(12844)

P=2L+2W
290=2(W+25)+2W
290=2W+50+2W
290=4W+50
4W=290-50
4W=240
W=240/4
W=60 YARDS FOR THE WIDTH.
L=60+25
L=85 YARDS FOR THE LENGTH.
PROOF:
290=2*85+2*60
290=170+120
290=290